# Training Latent Diffusion Models with Interacting Particle Algorithms

### v1

@@ Abstract
We introduce a novel particle-based algorithm for end-to-end training of latent diffusion models. We reformulate the training task as minimizing a free energy functional and obtain a gradient flow that does so. By approximating the latter with a system of interacting particles, we obtain the algorithm, which we underpin it theoretically by providing error guarantees. The novel algorithm compares favorably in experiments with previous particle-based methods and variational inference analogues.
@@ 3.3 A practical algorithm: IPLD
When training simpler latent variable models with algorithms similar to ours, Kuntz et al., (2023); Lim et al., (2024) observed that preconditioning the models’ parameters similarly as in RMSProp (Tieleman and Hinton,, 2012) mitigated ill-conditioning and stabilized the training. Lim et al., (2024) further noted gains in both training speeds and test-time performance by incorporating momentum into the parameter and particle updates. Here we precondition and incorporate momentum for both the parameters and particles. We use optimizers like AdamW (Loshchilov and Hutter,, 2019) for the former and adaptive Langevin algorithms (Li et al.,, 2016; Kim et al.,, 2020) for the latter. Lastly, we choose different step sizes $h_{\theta}$, $h_{\phi}$, and $h_{z}$ respectively for $\theta$, $\phi$, and the particles to account for the multi-scale nature of the interacting particle system (Akyildiz et al.,, 2024; Pavliotis and Stuart,, 2008).
@@
where $\gamma_{t}:[0,\infty)\to(0,\infty)$ denotes the (non-decreasing) annealing schedule. This is a practice commonly used when training deep Latent Variable Models (LVMs) to encourage accurate reconstruction during the early stages of training (Sønderby et al.,, 2016; Fu et al.,, 2019; Vahdat and Kautz,, 2020; Vahdat et al.,, 2021). Adjusting $\tilde{F}$ correspondingly results in $\gamma_{t}$ pre-multiplying $\mathcal{L}_{\mathcal{D}}$ in (10); which in turn corresponds to adjusting the noise levels in (14,17) by replacing $\sqrt{2h}$ therein with $\sqrt{2h\gamma_{t}}$ (cf. Section A.4.1 for details). We provide the pseudocode for the full training algorithm in Section A.4.
@@ 4 Related Work
There have been many attempts at incorporating DMs into LVMs as the prior $p_{\theta}(z)$. Vahdat et al., (2021) proposed to jointly train a continuous-time score-based DM (Song et al., 2021c, ) in the latent space of a deep hierarchical VAE (Vahdat and Kautz,, 2020). Similarly, Wehenkel and Louppe, (2021) considered the joint training of a discrete-time DM with a conventional VAE. Cohen et al., (2022) instead applied a discrete-time diffusion prior to a Vector-Quantised Variational Autoencoder (VQ-VAE) (van den Oord et al.,, 2017). The seminal work by Rombach et al., (2022) can also be viewed as learning a diffusion prior post training of the VAE. Concurrent to our work, Leng et al., (2025) also considers the joint training of a latent diffusion model with the VAE, but their objective is based instead on the REPresentation Alignment (REPA) loss (Yu et al.,, 2025). Additionally, several studies have proposed a trainable forward process for diffusion models (Bartosh et al.,, 2024; Nielsen et al.,, 2024; Kim et al.,, 2022). These models come with an additional inference network and operate in a latent space whose dimension equals that of the pixel space. More recently, combinations of diffusion-based priors with other types of probabilistic models in hierarchical VAEs have been explored, such as Energy-based Model (EBM) (Cui and Han,, 2024) and Variational Mixture of Posteriors prior (Kuzina and Tomczak,, 2024). Additionally, the work by Silvestri et al., (2025) can also be viewed as an VAE with a consistency model (Song et al.,, 2023) learning the aggregated posterior.
@@ 6 Discussion
Appendix A Derivation of the gradient flow
@@ A.1 Derivation of the Training Objective
We derive the training objective $\tilde{F}(\theta,\phi,q^{1:M})$ in (4) and include a derivation of the standard reparametrized diffusion objective $\mathcal{L}_{D}(\theta,z_{0})$ analogous to Ho et al., (2020); Sohl-Dickstein et al., (2015). For convenience, we re-state the objective here and drop the dependence on the $m$th data point:
@@
But note that we can re-write $q_{k}(z_{k}|z_{k-1})$ using Bayes’ rule as:
@@
$\mathcal{L}_{D}(\theta,z_{0})$ $=\int\left(\log\frac{q(z_{K}|z_{0})}{p_{K}(z_{K})}-\log{p_{\theta,1}(z_{0}|z_{1})}+\sum_{k=2}^{K}\log\frac{q_{k}(z_{k-1}|z_{k},z_{0})}{p_{\theta,k}(z_{k-1}|z_{k})}\right)q(z_{1:K}|z_{0}){\mathrm{d}}z_{1:K}.$ (22)
$=\mathbb{E}_{q(z_{1:K}|z_{0})}\left[D_{\mathrm{KL}}(q(z_{K}|z_{0})\|p(z_{K}))-\log p_{\theta,1}(z_{0}|z_{1})\right]$ (23)
@@
We set the transition kernels to be Gaussian and use the same parametrization as in Ho et al., (2020). For the forward process, we take:
@@
where $\{\beta_{k}\}_{k=1}^{K}$ is a linear noise schedule with $\beta_{k}=(1-k/K)\beta_{0}+(k/K)\beta_{K}$. Using the notation $\alpha_{k}:=1-\beta_{k}$ and $\bar{\alpha}_{k}:=\prod_{s=1}^{k}\alpha_{s}$, we can derive the $k$-step transition kernel:
@@
Note that in contrast to the standard pixel-space DDPM, we have the extra term in 33 depending on the latent $z_{0}$; we also use a Gaussian distribution for $p_{\theta}(z_{0}|z_{1})$ instead of a discrete one in 34.
@@
$D_{\mathrm{KL}}({q_{k}(z_{k-1}|z_{k},z_{0})}\|{p_{\theta,k}(z_{k-1}|z_{k})})$ $=\frac{1}{2}\log\left(\frac{1-\bar{\alpha}_{k}}{{\beta}_{k}}\right)-\frac{d_{z}}{2}+\frac{d_{z}{\beta}_{k}+\beta_{k}\|\epsilon-\epsilon_{\theta,k}(z_{k})\|^{2}}{2\alpha_{k}(1-\bar{\alpha}_{k})}$ (35)
@@
More specifically, we draw a minibatch of indices $\mathcal{B}\subseteq[M]$, standard Gaussian random variables $\epsilon^{m,n}\sim\mathcal{N}(0,I)$, and times $t^{m,n}\sim\text{Unif}(1,\ldots,K)$ for $m\in\mathcal{B},n\in[N]$ to approximate the loss by:
@@
A.2 Otto Calculus on $\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{\phi}}\times\mathcal{P}(\mathbb{R}^{d_{z}})$
To obtain the gradient flow in Section 3.1, we view the product space of parameters and probability distributions $\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{\phi}}\times\mathcal{P}(\mathbb{R}^{d_{z}})$ as a Riemannian manifold. We will equip this product space with suitable tangent spaces and Riemannian metrics, which enable us to define the gradients and perform optimization. We omit the subscript on $z_{0}$ and denote it by $z$ for simplicity.
@@
We first note that it is equivalent to derive the gradient flow on $\mathbb{R}^{D}\times\mathcal{P}(\mathbb{R}^{d_{z}})$, where $D=d_{\theta}+d_{\phi}$ is the sum of dimensions of the parameter spaces $\Phi$ and $\Theta$. Using the same set-up as in Kuntz et al., (2023), we denote the parameter space as $\Theta=\mathbb{R}^{D}$, and we assume the approximate posterior is a distribution with strictly positive densities w.r.t. the Lebesgue measure and have support ${\mathsf{Z}}=\mathbb{R}^{d_{z}}$. We can thus define the product manifold as $M=\Theta\times\mathcal{P}({\mathsf{Z}})$. For each point $(\theta,q)\in M$, we can define the tangent space $TM$ and its dual $T^{*}M$ as:
@@
where we note that $T_{\theta}\Theta\cong T_{\theta}^{*}\Theta\cong\mathbb{R}^{d_{\theta}}$. We also define the tangent and cotangent space of ${\mathsf{Z}}$ at $q$ as in (Otto,, 2001; Ambrosio et al.,, 2021):
@@
$T_{q}^{*}\mathcal{P}(\mathsf{Z})$ :=$:=\left\{f:\mathbb{R}^{d_{z}}\to\mathbb{R}\right\}/\mathbb{R}$
@@
Furthermore, we define the duality pairing, which is a triplet $(T_{q}\mathcal{P}({\mathsf{Z}}),T_{q}^{*}\mathcal{P}({\mathsf{Z}}),\langle\cdot,\cdot\rangle)$ with $\langle\cdot,\cdot\rangle:T_{q}\mathcal{P}({\mathsf{Z}})\times T_{q}^{*}\mathcal{P}({\mathsf{Z}})\to\mathbb{R}$ by:
@@
where $G_{(\theta,q)}:TM\to T^{*}M$ is an invertible, self-adjoint, and positive-definite linear map. We will only consider tensors in a ’block-diagonal’ form (cf. Lee, (2018, Chapter 2)), which defines the metric through:
@@
To perform gradient descent on manifolds, we need an analogue of the gradient for a smooth function $F:M\to\mathbb{R}$ as in the Euclidean space. This is a vector field $\nabla F:M\to TM$ satisfying:
$g_{(\theta,q)}(\nabla F(\theta,q),(\tau,m))=dF_{(\theta,q)}(\tau,m),\quad\forall(\tau,m)\in TM,\ (\theta,q)\in M,$ (46)
where $dF_{(\theta,q)}:TM\to\mathbb{R}$ is an element of the cotangent space $T^{*}M$. By expressing in local basis, we can compute the gradient as:
@@ A.3 Derivation of the Gradient Flow
Equipped with the tools above, we now give a derivation of the gradients with respect to the distribution $q$ for the free energy functional $\tilde{F}(\theta,\phi,q)$ in (19). We first need to compute the first variation of the functional:
@@
Given the free energy of the form in eq. 20:
@@
$\delta_{q}\tilde{F}\theta,\phi,q)$ $=\log\left(\frac{q(z)}{p_{\phi}(x|z)}\right)+\mathcal{L}_{D}(\theta,z)$ (50)
@@
For Equations 51 and 52, it suffices to note that by grouping the parameters into $\tilde{\theta}:=(\theta,\phi)\in\Theta=\mathbb{R}^{D}$, we can compute by Taylor expansion:
@@
For 50, we recall the first variation is linear, thus it suffices to compute $\delta_{q}\int\mathcal{L}_{D}(\theta,z)q(z){\mathrm{d}}z$ and $\delta_{q}\int\log\left(\frac{q(z)}{p_{\phi}(x|z)}\right)q(z){\mathrm{d}}z$ separately. Now note that:
@@
and for the other one we note that $\log(z+t)(z+t)=\log(z)z+[\log(z)+1]t+o(t)$, from which it follows:
$\int\log\left(\frac{q(z)+tm(z)}{p_{\phi}(x|z)}\right)[q(z)+tm(z)]{\mathrm{d}}z$ $=\int\log(q(z)+tm(z))[q(z)+tm(z)]{\mathrm{d}}z$
$-\int\log p_{\phi}(x|z)[q(z)+tm(z)]{\mathrm{d}}z$
=$=\int\left\{(\log q(z))q(z)+[\log q(z)+1]tm(z)+o(t)\right\}{\mathrm{d}}z$
@@
$=\int\log\left(\frac{q(z)}{p_{\phi}(x|z)}\right)q(z){\mathrm{d}}z+t\int\log\left(\frac{q(z)}{p_{\phi}(x|z)}\right)m(z){\mathrm{d}}z+o(t)$
where we have used that $\int m(z){\mathrm{d}}z=0$ as $m\in T\mathcal{P}(\mathsf{Z})$. ∎
@@
$\nabla_{\theta}\tilde{F}(\theta,\phi,q)$ $=\int\left[\nabla_{\theta}\mathcal{L}_{D}(\theta,z_{0})\right]q(z){\mathrm{d}}z$ (53)
@@
Recall that on the manifold $M=\Theta\times\mathcal{P}(\mathsf{Z})$, the gradient can be computed from the first variation via the metric tensor:
$\nabla_{(\theta,q)}\tilde{F}(\theta,q)=G_{(\theta,q)}^{-1}\delta\tilde{F}(\theta,q)=((G_{\theta})^{-1}\delta_{\theta}\tilde{F}(\theta,q),(\mathbf{G}_{q}^{W})^{-1}\delta_{q}\tilde{F}(\theta,q))$
Now note that for the parameters $\theta,\phi$, the gradient is equal to the first variation by considering $\theta=(\theta,\phi)$. For the distribution $q$, we recall the Wasserstein-2 metric tensor $(\mathbf{G}_{q}^{W})^{-1}f=-\nabla_{z}\cdot(q\nabla_{z}f)$ for all $f\in C^{\infty}(\mathcal{P}(\mathsf{Z}))$, which gives the desired result using Section A.3. ∎

### v2

@@ Abstract
We introduce a novel particle-based algorithm for end-to-end training of latent diffusion models. We reformulate the training task as minimizing a free energy functional and obtain a gradient flow that does so. By approximating the latter with a system of interacting particles, we obtain the algorithm, which we underpin theoretically by providing error guarantees. The novel algorithm compares favorably in experiments with previous particle-based methods and variational inference analogues.
@@ 3.3 A practical algorithm: IPLD
When training simpler latent variable models with algorithms similar to ours, Kuntz et al., (2023); Lim et al., (2024) observed that preconditioning the models’ parameters similarly as in RMSProp (Tieleman and Hinton,, 2012) mitigated ill-conditioning and stabilized the training. Lim et al., (2024) further noted gains in both training speeds and test-time performance by incorporating momentum into the parameter and particle updates. Here we precondition and incorporate momentum for both the parameters and particles. We use optimizers like AdamW (Loshchilov and Hutter,, 2019) for the former and adaptive Langevin algorithms (Li et al.,, 2016; Kim et al.,, 2020) for the latter. Lastly, we choose different step sizes $h_{\theta}$, $h_{\phi}$, and $h_{z}$ respectively for $\theta$, $\phi$, and the particles to account for the multi-scale nature of the interacting particle system (Akyildiz et al.,, 2024; Pavliotis and Stuart,, 2008). We provide the pseudocode for the full training algorithm in Section A.4.
@@
where $\gamma_{t}:[0,\infty)\to(0,\infty)$ denotes the (non-decreasing) annealing schedule. This is a practice commonly used when training deep Latent Variable Models (LVMs) to encourage accurate reconstruction during the early stages of training (Sønderby et al.,, 2016; Fu et al.,, 2019; Vahdat and Kautz,, 2020; Vahdat et al.,, 2021). Adjusting $\tilde{F}$ correspondingly results in $\gamma_{t}$ pre-multiplying $\mathcal{L}_{\mathcal{D}}$ in (10); which in turn corresponds to adjusting the noise levels in (14,17) by replacing $\sqrt{2h}$ therein with $\sqrt{2h\gamma_{t}}$ (cf. Section A.4.1 for details).
@@ 4 Related Work
There have been many attempts at incorporating DMs into LVMs as the prior $p_{\theta}(z)$. Vahdat et al., (2021) proposed to jointly train a continuous-time score-based DM (Song et al., 2021c, ) in the latent space of a deep hierarchical VAE (Vahdat and Kautz,, 2020). Similarly, Wehenkel and Louppe, (2021) considered the joint training of a discrete-time DM with a conventional VAE. Cohen et al., (2022) instead applied a discrete-time diffusion prior to a Vector-Quantised Variational Autoencoder (VQ-VAE) (van den Oord et al.,, 2017). The seminal work by Rombach et al., (2022) can also be viewed as learning a diffusion prior post training of the VAE. Concurrent to our work, Leng et al., (2025) also considers the joint training of a latent diffusion model with the VAE, but their objective is based instead on the REPresentation Alignment (REPA) loss (Yu et al.,, 2025). Additionally, several studies have proposed a trainable forward process for diffusion models (Kim et al.,, 2022; Bartosh et al., 2024a, ; Bartosh et al., 2024b, ; Nielsen et al.,, 2024), which can be viewed as generalizations of latent diffusion models. More recently, combinations of diffusion-based priors with other types of probabilistic models in hierarchical VAEs have been explored, such as Energy-based Model (EBM) (Cui and Han,, 2024) and Variational Mixture of Posteriors prior (Kuzina and Tomczak,, 2024). Additionally, the work by Silvestri et al., (2025) can also be viewed as an VAE with a consistency model (Song et al.,, 2023) learning the aggregated posterior.
@@ 6 Discussion
Appendix A Derivations
@@ A.1 Derivation of the Training Objective
We derive the training objective $\tilde{F}(\theta,\phi,q^{1:M})$ in (4) and include a derivation of the standard reparametrized diffusion objective $\mathcal{L}_{D}(\theta,z_{0})$ analogous to Ho et al., (2020); Sohl-Dickstein et al., (2015). For convenience, we re-state the objective for a single data point here and omit the superscript $m$:
@@
We can re-write $q_{k}(z_{k}|z_{k-1})$ using Bayes’ rule as:
@@
$\mathcal{L}_{D}(\theta,z_{0})$ $=\mathbb{E}_{q(z_{1:K}|z_{0})}\left[D_{\mathrm{KL}}(q(z_{K}|z_{0})\|p(z_{K}))-\log p_{\theta,1}(z_{0}|z_{1})\right]$ (22)
@@
For the forward process, we take:
@@
where $\{\beta_{k}\}_{k=1}^{K}$ is a linear noise schedule with $\beta_{k}=(1-k/K)\beta_{0}+(k/K)\beta_{K}$. Using the notation $\alpha_{k}:=1-\beta_{k}$ and $\bar{\alpha}_{k}:=\prod_{j=1}^{k}\alpha_{j}$, we can derive the $k$-step transition kernel:
@@
Note that in contrast to the standard pixel-space DDPM, we have the extra term in (32) depending on the latent $z_{0}$; we also use a Gaussian distribution with identity variance for $p_{\theta}(z_{0}|z_{1})$ in (33) instead of the independent discrete decoder from Ho et al., (2020).
Reparametrization.
@@
$D_{\mathrm{KL}}({q_{k}(z_{k-1}|z_{k},z_{0})}\|{p_{\theta,k}(z_{k-1}|z_{k})})$ $=\frac{1}{2}\log\left(\frac{1-\bar{\alpha}_{k}}{{\beta}_{k}}\right)-\frac{d_{z}}{2}+\frac{d_{z}{\beta}_{k}+\beta_{k}\|\epsilon-\epsilon_{\theta,k}(z_{k})\|^{2}}{2\alpha_{k}(1-\bar{\alpha}_{k})},$ (34)
@@
More specifically, we draw a minibatch of indices $\mathcal{B}\subseteq[M]$, standard Gaussian random variables $\epsilon^{m,n}\sim\mathcal{N}(0,I)$, and $t^{m,n}\sim\text{Unif}(1,\ldots,K)$ for $m\in\mathcal{B},n\in[N]$ to approximate the loss by:
@@
To simplify the notation, we will denote the one-sample approximation ^$\hat{\mathcal{L}}_{\text{simple}}(\theta,z_{0}^{m,n};\epsilon^{m,n},t^{m,n})$ with ^$\hat{\mathcal{L}}_{\text{simple}}(\theta,z_{0}^{m,n})$ as in Section 3.3 and we use the same simplified notation in the rest of the appendix.
A.2 The Euclidean-Wasserstein Geometry on $\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{\phi}}\times\mathcal{P}(\mathbb{R}^{d_{z}})$
To obtain the gradient flow in Section 3.1, we view the product space of parameters and probability distributions $\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{\phi}}\times\mathcal{P}(\mathbb{R}^{d_{z}})$ as a Riemannian manifold. We will equip this product space with suitable tangent spaces and Riemannian metrics, which enable us to define the gradients and perform optimization. We omit the subscript on $z_{0}$ and denote it by $z$ for simplicity throughout this section.
@@
We first note that it is equivalent to derive the gradient flow on the combined $\mathbb{R}^{D}\times\mathcal{P}(\mathbb{R}^{d_{z}})$, where $D=d_{\theta}+d_{\phi}$ is the sum of dimensions of the parameter spaces. Using the same set-up as in Kuntz et al., (2023), we denote the parameter space as $\Theta=\mathbb{R}^{D}$, and we assume the approximate posterior is a distribution with strictly positive densities w.r.t. the Lebesgue measure and have support ${\mathsf{Z}}=\mathbb{R}^{d_{z}}$. We let the product manifold be $M:=\Theta\times\mathcal{P}({\mathsf{Z}})$. For each point $(\theta,q)\in M$, we can define the tangent space $TM$ and its dual $T^{*}M$ as:
@@
where we note that $T_{\theta}\Theta\cong T_{\theta}^{*}\Theta\cong\mathbb{R}^{D}$. We also define the tangent and cotangent space of $\mathcal{P}({\mathsf{Z}})$ at $q$ as in Otto, (2001); Ambrosio et al., (2021):
@@
$T_{q}^{*}\mathcal{P}(\mathsf{Z})$ :=$:=\left\{f:{\mathsf{Z}}\to\mathbb{R}\right\}/\mathbb{R}$
@@
Furthermore, we define the duality pairing, which is a triplet $(T_{q}\mathcal{P}({\mathsf{Z}}),T_{q}^{*}\mathcal{P}({\mathsf{Z}}),\langle\cdot,\cdot\rangle)$ with $\langle\cdot,\cdot\rangle:T_{q}\mathcal{P}({\mathsf{Z}})\times T_{q}^{*}\mathcal{P}({\mathsf{Z}})\to\mathbb{R}$ given by:
@@
where $G_{(\theta,q)}:TM\to T^{*}M$ is an invertible, self-adjoint, and positive-definite linear map. We will only consider tensors in a ’block-diagonal’ form (cf. Lee, (2018, Chapter 2)), which defines the metric via:
@@
To perform gradient descent on manifolds, we further need an analogue of the gradient for a smooth function $F:M\to\mathbb{R}$ as in the Euclidean space. This is a vector field $\nabla F:M\to TM$ satisfying:
$g_{(\theta,q)}(\nabla F(\theta,q),(\tau,m))=\lim_{t\to 0}\frac{F(\theta+t\tau,q+tm)-F(\theta,q)}{t}\quad\forall(\tau,m)\in TM,\;(\theta,q)\in M.$ (45)
By expressing in local coordinates, we can compute the gradient as:
@@ A.3 Derivation of the Gradient Flow
Equipped with the tools above, we now give a derivation of the gradients with respect to the distribution $q$ for the free energy functional $\tilde{F}(\theta,\phi,q)$ in (19). We first need to compute the first variation (47) of the functional:
@@
Given the free energy of the form in (20):
@@
$\delta_{q}\tilde{F}(\theta,\phi,q)$ $=\log\left(\frac{q(z)}{p_{\phi}(x|z)}\right)+\mathcal{L}_{D}(\theta,z)$ (49)
@@
For eqs. 50 and 51, it suffices to note that by grouping the parameters into $\tilde{\theta}:=(\theta,\phi)\in\Theta=\mathbb{R}^{D}$, we can compute by Taylor expansion:
@@
For (49), we recall the first variation is linear, thus it suffices to compute $\delta_{q}\int\mathcal{L}_{D}(\theta,z)q(z){\mathrm{d}}z$ and $\delta_{q}\int\log\left(\frac{q(z)}{p_{\phi}(x|z)}\right)q(z){\mathrm{d}}z$ separately. Now note that:
@@
and for the latter we note that $\log(z+t)(z+t)=\log(z)z+[\log(z)+1]t+o(t)$, from which it follows:
$\int\log\left(\frac{q(z)+tm(z)}{p_{\phi}(x|z)}\right)[q(z)+tm(z)]{\mathrm{d}}z$ $=\int(\log q(z))q(z)+[\log q(z)+1]tm(z)+o(t){\mathrm{d}}z$
@@
$=\int\log\left(\frac{q(z)}{p_{\phi}(x|z)}\right)q(z){\mathrm{d}}z$
$+t\int\log\left(\frac{q(z)}{p_{\phi}(x|z)}\right)m(z){\mathrm{d}}z+o(t)$
where we have used that $\int m(z){\mathrm{d}}z=0$ for all $m\in T\mathcal{P}(\mathsf{Z})$. ∎
@@
$\nabla_{\theta}\tilde{F}(\theta,\phi,q)$ $=\int\left[\nabla_{\theta}\mathcal{L}_{D}(\theta,z)\right]q(z){\mathrm{d}}z$ (52)
@@
Recall that on the manifold $M=\Theta\times\mathcal{P}(\mathsf{Z})$, the gradient can be computed from the first variation via the metric tensor (46):
$\nabla_{(\tilde{\theta},q)}\tilde{F}(\tilde{\theta},q)=G_{(\tilde{\theta},q)}^{-1}\delta\tilde{F}(\tilde{\theta},q)=((G_{\theta})^{-1}\delta_{\theta}\tilde{F}(\tilde{\theta},q),(\mathbf{G}_{q}^{W})^{-1}\delta_{q}\tilde{F}(\tilde{\theta},q)),$
where $\tilde{\theta}=(\theta,\phi)$. For the parameters $\theta,\phi$, their gradients are given by the usual Euclidean gradients. For the distribution $q$, we recall the Wasserstein-2 metric tensor $(\mathbf{G}_{q}^{W})^{-1}f=-\nabla_{z}\cdot(q\nabla_{z}f)$ for all $f\in C^{\infty}(\mathcal{P}(\mathsf{Z}))$, which gives the desired result using Section A.3. ∎